# DistALANER: Distantly Supervised Active Learning Augmented Named Entity Recognition in the Open Source Software Ecosystem

### 25 Feb 2024 v1

@@ Abstract
This paper proposes a novel named entity recognition (NER) technique specifically tailored for the open-source software systems. Our approach aims to address the scarcity of annotated software data by employing a comprehensive two-step distantly supervised annotation process. This process strategically leverages language heuristics, unique lookup tables, external knowledge sources, and an active learning approach. By harnessing these powerful techniques, we not only enhance model performance but also effectively mitigate the limitations associated with cost and the scarcity of expert annotators. It is noteworthy that our framework significantly outperforms the state-of-the-art LLMs by a substantial margin. We also show the effectiveness of NER in the downstream task of relation extraction. To ensure transparency and facilitate future research in this domain, we are committed to making all our codes, datasets, and models openly available in the public domain upon acceptance [^1] https://anonymous.4open.science/r/DistALANER-DC9D
/README.md
DistALANER: Distantly Supervised Active Learning Augmented Named Entity Recognition in the Open Source Software Ecosystem
[…]
@@ 1 Introduction
Traditional named entity recognition (NER) models exhibit certain limitations, particularly when dealing with domain-specific data. Primarily, this stems from the fact that NER models are conventionally trained on generic corpora, rendering them less effective when encountering text sourced from specialized domains such as software, legal Trias et al. (2021), biomedical Jensen et al. (2006); Krallinger et al. (2015); Nedellec et al. (2013) or engineering fields Li et al. (2022), which inherently possess distinctive vocabularies and entities. For instance, the word “windows” could denote a well-known operating system in the realm of software and technology, yet simultaneously refer to a commonplace architectural feature in the context of residential or commercial structures. Adapting a generic NER model to handle tasks specific to a certain field often requires extra, specialized training data from that area. However, gathering and annotating this additional data can be a costly and time-consuming task.
@@
In the area of open source software development, the need for NER has become increasingly critical Tang et al. (2023). NER plays a pivotal role in deciphering and categorizing textual information into predefined entities such as individual contributors, programming languages, software tools, and project specifications found in software documentation, source code, bug reports, or community discussions Wang et al. (2021). The understanding and categorization of such entities offer deep insights, allowing for effective communication, resource allocation, and decision-making within the open source community. In addition, it assists in knowledge extraction Perera et al. (2020), facilitating a greater comprehension of the system’s functionality Shrimal et al. (2022), feature requirements, software dependencies, or potential bugs Zhou et al. (2020), thereby augmenting the efficiency of software development and maintenance. Further, NER can aid in community management tasks, such as identifying contributors and their areas of expertise or mapping the interactions within the developer community. Consequently, the integration of NER in the open source software domain could dramatically streamline processes, enhance collaboration, and eventually improve the overall quality of the software produced.
@@
Recently, Large Language Models (LLMs) can be highly effective in the software domain Ozkaya (2023). With their ability to understand complex patterns and generate human-like text, they can assist in identifying and classifying key entities such as specific coding languages, software tools, packages, peripherals, or developers mentioned in various sources like code, software documentation, and community discussions. However, LLMs can sometimes prove to be a bottleneck in identifying NER due to security issues, cost and lack of contextual knowledge. These limitations arise due to their broad contextual learning from vast corpora that often spans numerous domains, making them less specialized for a particular field like software development. These models might struggle to identify and classify domain-specific entities accurately, given their generic training Qin et al. (2023). Moreover, as LLMs learn from data available up to their last training checkpoint, they may not be aware of new terms or entities introduced in the domain post-training. Taking into account the aforementioned limitations Yang et al. (2023), we put forth an innovative framework, explicitly designed for the software domain trained through bug data, manuals and CQAs. Our objective is to incorporate lightweight models into our framework, allowing them to seamlessly integrate with any system. Our approach involves limited human intervention. By including a variety of methods depicted in Figure 2, we enhance the efficiency of deep learning models. The key contributions of this paper are as follows.
@@
  • In terms of datasets we release the following items - (a) a huge open source domain corpus with both human and system annotations for nine different named entity types (see examples in Figure 1), (b) For each entity type, a large unique lookup table contains relevant entities automatically collected from various sources since 2004 (to check detailed sources, see appendix) and (c) a huge corpus of human annotated entity relation pairs in software domain for the downstream application task [^2] https://zenodo.org/record/8075578.
@@ 2 Related work
Named entity recognition in software ecosystems: Initially Ye et al. (2016) proposed S-NER for identifying software-related entities, focusing on high-level entity types such as programming language and API, amongst others. However, they purely rely on human annotations for their model training and testing. Li et al. (2020); Zhou et al. (2020) introduced a unified framework employing machine reading comprehension for nested and flat NER. Tabassum et al. (2020) presented SoftNER for recognizing code tokens and software-related entities in natural language and programming texts.
Distant supervision based named entity recognition: In Wang et al. (2019); Fang et al. (2021) the authors introduced a novel method of dictionary extension for extracting new entities utilizing a type expanded model. Another work by Yu Meng ; Peng et al. (2019) presented a novel method for training NER models using distantly labeled data, typically acquired by associating entity mentions in raw text with corresponding entity types. Liang et al. (2020) put forward a unique computational framework, BOND, which leverages pretrained language models to enhance NER model performance. GPT-NER by Wang et al. (2023) is a first of its kind approach that recasts the sequence labeling task into a generative framework.
Our work introduces a distantly supervised method to auto-annotate named entities in the open-source software systems. This approach is very effective is a low resource setting where gold labels are scarce. We further show that this NER pipeline improves the task of relation extraction.
@@ 3 Dataset
Ubuntu bug repository: In our experiment, we use the repository of bugs collected by (Hazra et al., 2021a). These bugs are mainly reported on packages, conflicts between Ubuntu and Windows and other related events. The dataset contains approximately 270K bugs along with the metadata such as title of the bug, description of the bug, user who posted the bug, comments and their commenters, creation date of the bug, tags of the bugs. In our work, we mainly use the description of the bugs to obtain the named entities. We filter this raw dataset by excluding (i) those bugs that solely reference another bug, for example, “Automatically imported from Debian bug report #257568" [^6] http://bugs.debian.org/257568 and (ii) those bugs that have very small description size ($<60$ words) or exceedingly large description size ($>400$ words). The bugs with very small description size prohibits obtaining meaningful representations while those with very large size routinely include code logs in massive proportions rather than useful text. The dataset [^7] https://launchpad.net/ statistics are presented in Table 1.
@@ 5.1 Stage 1: Construction and matching of dictionary
At the first stage, we build the dictionary of entities and their respective entity types. To build the dictionary, we use existing knowledge from websites, repositories, and documents of the Ubuntu eco system. For the OS entity, we include all the Ubuntu distributions (obtained from (Hazra et al., 2021a)) and collect Linux distributions and Windows versions from Wikipedia. For the ARC entity type, we include different writing styles (and versions) of 32 bit and 64 bit (x32 | x64, x386 | amd64, etc.). For CMD type, we collect all the Linux commands from Wikipedia, Ubuntu man pages, and the paper (Zhang et al., 2022). Authors in (Hazra et al., 2021a) collected the list of packages of each Ubuntu distributions (total 20 distributions). For the PKG type, we utilize this list and extract all the unique package names. We build the ERR type by collecting the error codes that usually occur in the Ubuntu system from the Ubuntu wiki page [^8] https://wiki.ubuntu.com/error/and/warning/messages. For the type EXT, we collect data from various Wikipedia pages. For the ORG type, we focus on software organizations and collect the information from Wikipedia. For type PRP, we gather entities from various sources (see Appendix C) followed by manual inspection. For SOC, we collect data from various unix-related websites as well as Wikipedia pages (See Table 2). Next, we conduct dictionary matching. We consider only exact string matching of the bug description text with the entities corresponding to each entity type. For instance, ‘CurrentDesktop’ is not marked because it contains ‘Desktop’ as a subword. We discard the wrongly identified entities using a set of regex (see Appendix D). Note that, in earlier papers Shang et al. (2018); Wang et al. (2020), researchers have used the AutoPhrase tool  [^9] https://github.com/shangjingbo1226/AutoPhrase to extract phrases from the text, which they later consider for entity identification. We also attempted to apply AutoPhrase to identify entity-containing phrases. However, the tool failed to extract key phrases from our bug descriptions and instead assigns higher scores to irrelevant terms like ‘box’, ‘release’, ‘start’, ‘architecture’, ‘show’, ‘network’, ‘subdevice’, and ‘software’. Hence, we resorted to the above dictionary matching technique.
@@ 5.2 Stage 2: Entity distillation and dictionary expansion
In Stage 2, we distill the exactly matched entities and then expand the list of entities for a given entity type using an active learning approach. Active learning enables our system to learn iteratively, refining its understanding as it receives feedback from its interactions. We observe that many phrases get marked as an entity, but in reality, some of them are not entities. Thus, we employ two heuristics to distill the entities: (1) identifying parts-of-speech patterns, and (2) human intervention. Parts-of-speech patterns: After performing exact matching, we find that the likelihood of a phrase being an entity often depends on the parts-of-speech tag of the words in the phrase or before and after the phrase. For instance, the verb ‘find’ in a sentence is never an entity, even though our dictionary matching may mistakenly identify it as a CMD entity (see more samples in Appendix A). We adjust the contents of the dictionary through such iterative revisions. Human intervention: In this step, we avail human intervention to identify cases where a phrase has been incorrectly tagged as an entity. Active learning is particularly useful here, as the model can learn from the feedback provided by the human expert, improving its future predictions. We randomly sample a few automatically annotated bugs from all the bugs and check which phrases could potentially be non-entities. We encounter a challenge as the number of entities for categories like software components and peripherals is quite low in our entity list. To expand this dictionary, we utilize the software tool, TagMe [^10] https://sobigdata.d4science.org/web/tagme/tagme-help, to extract mentions from Wikipedia and consider them as entities. Along with that process, after receiving the annotated data from Stage 1 and Parts-of-speech filtered bugs (L) also with their IO mentions, a binary RoBERTa-based classifier is trained. We sample 100 bugs on a yearly basis, with our sample spanning from 2004 to 2019. Subsequently, we take up 100*16 = 1600 bugs from TagMe but not included in Stage 1 and feed them into our model for classification as either “entity” or “non-entity”. If the confidence for an entity being identified is 50% or greater, it undergoes manual labeling by a human to categorize it into one of the nine predefined entity types. Once the manual annotation is completed, the newly labeled data (d) is integrated into L, resulting in an updated dataset, (L $\cup$ d). The model is then retrained using this revised dataset, and the entire procedure is reiterated until the stopping criterion is met. The stopping criterion is defined as the absence of data in the list of entities identified by TagMe. Finally, our requirement for human intervention end up with around $\sim$3682 filtered mentions, a relatively small number considering the size of our dataset which contains 170K bugs (equivalent to $\sim$1.2 million entities). Distributions of identified entities are shown in Figure 3.
@@ 5.3 Stage 3: The NER model
We use a variety of NER models for obtaining the final tags. These include Linear-CRF Lafferty et al. (2001), BilSTM-CRF Huang et al. (2015), BERT-CRF Souza et al. (2020), BERT-NER Liu et al. (2021), RoBERTa-CRF Jurkiewicz et al. (2020), SpanBERT-CRF Portelli et al. (2021), and SoftNER Tabassum et al. (2020). All models are implemented to execute task-specific actions, with their performance evaluated using precision, recall, and F1-score. However, recall is highlighted in this paper as a representational metric for NER tasks. We train the existing CRF based models using our distantly labelled data obtained from Stage 1 and Stage 2. We note the different hyperparameters for the above models in Appendix B. We also use LLMs in a zero-shot setting with instruction based prompt to extract the entities along with its start and end index from the paragraph (see Appendix I for more details).

### 15 Mar 2024 v3

@@ Abstract
With the AI revolution in place, the trend for building automated systems to support professionals in different domains such as the open source software systems, healthcare systems, banking systems, transportation systems and many others have become increasingly prominent. A crucial requirement in the automation of support tools for such systems is the early identification of named entities, which serves as a foundation for developing specialized functionalities. However, due to the specific nature of each domain, different technical terminologies and specialized languages, expert annotation of available data becomes expensive and challenging. In light of these challenges, this paper proposes a novel named entity recognition (NER) technique specifically tailored for the open-source software systems. Our approach aims to address the scarcity of annotated software data by employing a comprehensive two-step distantly supervised annotation process. This process strategically leverages language heuristics, unique lookup tables, external knowledge sources, and an active learning approach. By harnessing these powerful techniques, we not only enhance model performance but also effectively mitigate the limitations associated with cost and the scarcity of expert annotators. It is noteworthy that our model significantly outperforms the state-of-the-art LLMs by a substantial margin. We also show the effectiveness of NER in the downstream task of relation extraction  [^1] https://anonymous.4open.science/r/DistALANER-DC9D/README.md.
@@ 1 Introduction
Traditional named entity recognition (NER) models exhibit certain limitations, particularly when dealing with domain-specific data. Primarily, this stems from the fact that NER models are conventionally trained on generic corpora, rendering them less effective when encountering text sourced from specialized domains such as software, legal Trias et al. (2021), biomedical Jensen et al. (2006); Krallinger et al. (2015); Nédellec et al. (2013) or engineering fields Li et al. (2022), which inherently possess distinctive vocabularies and entities. For instance, the word “windows” could denote a well-known operating system in the realm of software and technology, yet simultaneously refer to a commonplace architectural feature in the context of residential or commercial structures. Adapting a generic NER model to handle tasks specific to a certain field often requires extra, specialized training data from that area. However, gathering and annotating this additional data can be a costly and time-consuming task.
@@
In the area of open source software development, the need for NER has become increasingly critical Tang et al. (2023). NER plays a pivotal role in deciphering and categorizing textual information into predefined entities such as individual contributors, programming languages, software tools, and project specifications found in software documentation, source code, bug reports, or community discussions Wang et al. (2022). The understanding and categorization of such entities offer deep insights, allowing for effective communication, resource allocation, and decision-making within the open source community. In addition, it assists in knowledge extraction Perera et al. (2020), facilitating a greater comprehension of the system’s functionality Shrimal et al. (2022), feature requirements, software dependencies, or potential bugs Zhou et al. , thereby augmenting the efficiency of software development and maintenance. Further, NER can aid in community management tasks, such as identifying contributors and their areas of expertise or mapping the interactions within the developer community. Consequently, the integration of NER in the open source software domain could dramatically streamline processes, enhance collaboration, and eventually improve the overall quality of the software produced.
@@
Recently, Large Language Models (LLMs) can be highly effective in the software domain Ozkaya (2023). With their ability to understand complex patterns and generate human-like text, they can assist in identifying and classifying key entities such as specific coding languages, software tools, packages, peripherals, or developers mentioned in various sources like code, software documentation, and community discussions. However, LLMs can sometimes prove to be a bottleneck in identifying NER due to security issues, cost and lack of contextual knowledge Banerjee et al. (2024). These limitations arise due to their broad contextual learning from vast corpora that often spans numerous domains, making them less specialized for a particular field like software development. These models might struggle to identify and classify domain-specific entities accurately, given their generic training Qin et al. (2023). Moreover, as LLMs learn from data available up to their last training checkpoint, they may not be aware of new terms or entities introduced in the domain post-training. Taking into account the aforementioned limitations Yang et al. (2023), we put forth an innovative framework, explicitly designed for the software domain trained through bug data, manuals and CQAs. Our objective is to incorporate lightweight models into our framework, allowing them to seamlessly integrate with any system. Our approach involves limited human intervention. By including a variety of methods depicted in Figure 2, we enhance the efficiency of deep learning models. The key contributions of this paper are as follows.
@@
  • In terms of datasets we release the following items - (a) a huge open source domain corpus with both human and system annotations for nine different named entity types (see examples in Figure 1), (b) For each entity type, a large unique lookup table contains relevant entities automatically collected from various sources since 2004 (to check detailed sources, see section 4) and (c) a huge corpus of human annotated entity relation pairs in software domain for the downstream application task [^2] https://zenodo.org/record/8075578.
@@ 2 Related work
Named entity recognition in software ecosystems: Initially Ye et al. (2016) proposed S-NER for identifying software-related entities, focusing on high-level entity types such as programming language and API, amongst others. They used a subset of the 2015 StackOverflow dataset and manually identified software-specific entities. Human annotators labeled additional posts, with the annotated data serving for supervised learning and model testing. Li et al. (2020); Zhou et al. introduced a unified framework employing machine reading comprehension for nested and flat Named Entity Recognition (NER). The authors devised a technique to extract the answer span for an entity, instead of labeling sequences. Extensive experimentation was carried out on nested and flat NER datasets to verify their proposed model. Tabassum et al. (2020) presented SoftNER for recognizing code tokens and software-related entities in natural language and programming texts. They used a programming-related dataset from StackOverflow, introducing an NER corpus for the programming domain. This corpus, annotated with 20 self-defined entity types, was used to train BERT Devlin et al. (2019), showing improvement over a pretrained model. In another study Kocaman and Talby (2020), the authors presented BiLSTM-CNN-Char, a deep learning architecture for extracting entities from biomedical notes and reports. The architecture was claimed to be efficient without relying on heavy transformer-based contextual embeddings. Recently, few-shot, multigrained, and nested NER have gained attention Ding et al. (2021). A notable contribution is the FEW-NERD dataset, a large-scale human-annotated collection featuring two-level entity type annotations. The authors in Xia et al. (2019) proposed Multi-Grained NER (MGNER), capable of identifying named entities at multiple granularities. The framework can handle non-overlapping or nested entity mentions and consists of a detector and a classifier component for efficient entity recognition.
Distant supervision based named entity recognition: In Wang et al. (2019); Fang et al. (2021) the authors introduced a novel method of dictionary extension for extracting new entities utilizing a type expanded model. This method’s primary aim was to enhance the performance of distantly supervised NER methods in automating data labeling and entity identification. The evaluation of their approach on diverse datasets demonstrated superior performance in comparison to the existing state-of-the-art distantly supervised systems. Another work by Yu Meng ; Peng et al. (2019) presented a novel method for training NER models using distantly labeled data, typically acquired by associating entity mentions in raw text with corresponding entity types. Liang et al. (2020) put forward a unique computational framework, BOND, which leverages pretrained language models to enhance NER model performance. GPT-NER by Wang et al. (2023) is a first of its kind approach that recasts the sequence labeling task into a generative framework.
Our work introduces a distantly supervised method to auto-annotate named entities in the open-source software systems. This approach is very effective is a low resource setting where gold labels are very limited. We further show that this NER pipeline improves the task of relation extraction.
@@ 3 Dataset
Ubuntu bug repository: In our experiment, we use the repository of bugs collected by Hazra et al. (2021). These bugs are mainly reported on packages, conflicts between Ubuntu and Windows and other related events. The dataset contains approximately 270K bugs along with the metadata such as title of the bug, description of the bug, user who posted the bug, comments and their commenters, creation date of the bug, tags of the bugs. In our work, we mainly use the description of the bugs to obtain the named entities. We filter this raw dataset by excluding (i) those bugs that solely reference another bug, for example, “Automatically imported from Debian bug report #257568" [^6] http://bugs.debian.org/257568 and (ii) those bugs that have very small description size ($<60$ words) or exceedingly large description size ($>400$ words). The bugs with very small description size prohibits obtaining meaningful representations while those with very large size routinely include code logs in massive proportions rather than useful text. The dataset [^7] https://launchpad.net/ statistics are presented in Table 1.
@@
4 Source details
  • Packages: We use all the packages assembled by the authors in Hazra et al. (2021).
  • Error codes: We collect these from the Ubuntu page [^12] https://wiki.ubuntu.com/error_and_warning_messages, and use TagMe (with our active learning based approach) to augment the list with additional error codes.
  • File extensions: We collect this data primarily from the Wikipedia page [^13] https://en.wikipedia.org/wiki/Filename_extension.
@@ 6.1 Stage 1: Construction and matching of dictionary
At the first stage, we build the dictionary of entities and their respective entity types. To build the dictionary, we use existing knowledge from websites, repositories, and documents of the Ubuntu eco system. For the OS entity, we include all the Ubuntu distributions (obtained from Hazra et al. (2021)) and collect Linux distributions and Windows versions from Wikipedia. For the ARC entity type, we include different writing styles (and versions) of 32 bit and 64 bit (x32 | x64, x386 | amd64, etc.). For CMD type, we collect all the Linux commands from Wikipedia, Ubuntu man pages, and the paper Zhang et al. (2022). Authors in Hazra et al. (2021) collected the list of packages of each Ubuntu distributions (total 20 distributions). For the PKG type, we utilize this list and extract all the unique package names. We build the ERR type by collecting the error codes that usually occur in the Ubuntu system from the Ubuntu wiki page [^14] https://wiki.ubuntu.com/error/and/warning/messages. For the type EXT, we collect data from various Wikipedia pages. For the ORG type, we focus on software organizations and collect the information from Wikipedia. For type PRP, we gather entities from various sources (see section 4 followed by manual inspection. For SOC, we collect data from various unix-related websites as well as Wikipedia pages (See Table 2). Next, we conduct dictionary matching. We consider only exact string matching of the bug description text with the entities corresponding to each entity type. For instance, ‘CurrentDesktop’ is not marked because it contains ‘Desktop’ as a subword. We discard the wrongly identified entities using a set of regex (see Appendix B). Note that, in earlier papers Shang et al. (2018); Wang et al. (2020), researchers have used the AutoPhrase tool  [^15] https://github.com/shangjingbo1226/AutoPhrase to extract phrases from the text, which they later consider for entity identification. We also attempted to apply AutoPhrase to identify entity-containing phrases. However, the tool failed to extract key phrases from our bug descriptions and instead assigns higher scores to irrelevant terms like ‘box’, ‘release’, ‘start’, ‘architecture’, ‘show’, ‘network’, ‘subdevice’, and ‘software’. Hence, we resorted to the above dictionary matching technique.
@@ 6.2 Stage 2: Entity distillation and dictionary expansion
In Stage 2, we distill the exactly matched entities and then expand the list of entities for a given entity type using an active learning approach. Active learning enables our system to learn iteratively, refining its understanding as it receives feedback from its interactions. We observe that many phrases get marked as an entity, but in reality, some of them are not entities. Thus, we employ two heuristics to distill the entities: (1) identifying parts-of-speech patterns, and (2) human intervention. Parts-of-speech patterns: After performing exact matching, we find that the likelihood of a phrase being an entity often depends on the parts-of-speech tag of the words in the phrase or before and after the phrase. For instance, the verb ‘find’ in a sentence is never an entity, even though our dictionary matching may mistakenly identify it as a CMD entity (see more samples in section 7). We adjust the contents of the dictionary through such iterative revisions. Human intervention: In this step, we avail human intervention to identify cases where a phrase has been incorrectly tagged as an entity. Active learning is particularly useful here, as the model can learn from the feedback provided by the human expert, improving its future predictions. We randomly sample a few automatically annotated bugs from all the bugs and check which phrases could potentially be non-entities. We encounter a challenge as the number of entities for categories like software components and peripherals is quite low in our entity list. To expand this dictionary, we utilize the software tool, TagMe [^16] https://sobigdata.d4science.org/web/tagme/tagme-help, to extract mentions from Wikipedia and consider them as entities. Along with that process, after receiving the annotated data from Stage 1 and Parts-of-speech filtered bugs (L) also with their IO mentions, a binary RoBERTa-based classifier is trained. We sample 100 bugs on a yearly basis, with our sample spanning from 2004 to 2019. Subsequently, we take up 100*16 = 1600 bugs from TagMe but not included in Stage 1 and feed them into our model for classification as either “entity” or “non-entity”. If the confidence for an entity being identified is 50% or greater, it undergoes manual labeling by a human to categorize it into one of the nine predefined entity types. Once the manual annotation is completed, the newly labeled data (d) is integrated into L, resulting in an updated dataset, (L $\cup$ d). The model is then retrained using this revised dataset, and the entire procedure is reiterated until the stopping criterion is met. The stopping criterion is defined as the absence of data in the list of entities identified by TagMe. Finally, our requirement for human intervention end up with around $\sim$3682 filtered mentions, a relatively small number considering the size of our dataset which contains 170K bugs (equivalent to $\sim$1.2 million entities). Distributions of identified entities are shown in Figure 3.
@@ 6.3 Stage 3: The NER model
We use a variety of NER models for obtaining the final tags. These include Linear-CRF Lafferty et al. (2001), BilSTM-CRF Huang et al. (2015), BERT-CRF Souza et al. (2020), BERT-NER Liu et al. (2021), RoBERTa-CRF Jurkiewicz et al. (2020), SpanBERT-CRF Portelli et al. (2021), and SoftNER Tabassum et al. (2020). All models are implemented to execute task-specific actions, with their performance evaluated using precision, recall, and F1-score. However, recall is highlighted in this paper as a representational metric for NER tasks. We train the existing CRF based models using our distantly labelled data obtained from Stage 1 and Stage 2. We note the different hyperparameters for the above models in Appendix D. We also use LLMs in a zero-shot setting with instruction based prompt to extract the entities along with its start and end index from the paragraph (see section 10 for more details).
7 Heuristics
Sample Bugs Sample Tag Heuristics Conversion
"Some selected error messages from the time of session login"
Messages - (NN, NNS, IN)
Time - (DT, NN, IN
Error ->O
Package - >O
rely on Xvfb with a less than optimal experience."
Less - (DT, JJR, IN) Package ->O
File - (DT, NN, IN)
File - (NNP, NN, IN)
File - (JJ, NN, IN)
Package ->O
Package ->O
Package ->O